satellite73 (satellite73):

the sequence is 2, 10, 30, 68,130, ... third differences is 6 so it is cubic what is the formula for each term?

OpenStudy (anonymous):

i got the answer from my eyeballs, but the question is: what is the method for deriving the sequence?

OpenStudy (anonymous):

what's the pattern in this sequence :/

OpenStudy (experimentx):

man ... are you trying to guess by looking? did you check my last answer ... my second last drawing!!

OpenStudy (zarkon):

\[n^3+n\]

OpenStudy (anonymous):

@satellite73 please tell what you got from your eyeballs

OpenStudy (anonymous):

sequence is \(k^3+k\) and yes, i guessed it, it wasn't hard you know it is cubic \[2=1^3+1\]\[10=2^3+2\]\[30=3^3+3\] pattern is clear, i just wanted a snap method for getting it

OpenStudy (anonymous):

npower3+n

OpenStudy (zarkon):

I used regression

OpenStudy (anonymous):

@Zarkon expand...

OpenStudy (anonymous):

please

OpenStudy (zarkon):

I found the cubic regression formula for the above 'data' set

OpenStudy (zarkon):

it is \[\hat{y}=x^3+x\]

OpenStudy (anonymous):

hmmm you mean a system of equations?

OpenStudy (zarkon):

A polynomial of degree 4 can fit in the above data...do a 'linear' regression with \[y=ax^4+bx^3+cx^2+dx+e\] you will get \[a=0,b=1,c=0,d=1,e=0\]

OpenStudy (zarkon):

let \[x=\{1,2,3,4,5\}\] and \[y=\{2, 10, 30, 68,130\}\]

OpenStudy (anonymous):

ok thanks. i think in this case maybe it was easier to guess, but in general not

OpenStudy (zarkon):

sure

OpenStudy (anonymous):

oh, but i know it is a polynomial of degree 3, because the third differences are constant. i can ignore the last term then? and use \(x=\{1,2,3,4\}\) \(y=\{2,10,30,68\}\)

OpenStudy (zarkon):

sure

OpenStudy (anonymous):

thnx

OpenStudy (zarkon):

NP

OpenStudy (experimentx):

Here is another one \[ 2 + \sum_{i = 1}^{n-1}\left( 8 + \sum_{j=1}^{i-1} 12 + (j-1)6\right)\] Mathematica code .. 2 + Sum[8 + Sum[12 + (i - 1) 6, {i, 1, j - 1}], {j, 1, n - 1}] though W|A doesn't like it